If the sum of three times a number and -5 is multiplied by 6, the answer is 42. Find the number.
Found 2 solutions by algebrahouse.com, ikleyn:
Answer by algebrahouse.com(1659)

x = the number

6[3x + (-5)] = 42 {sum of three times a number and -5 is multiplied by 6, answer is 42}
6(3x - 5) = 42 {simplified signs in brackets}
3x - 5 = 7 {divided each side by 6}
3x = 12 {added 5 to each side}
x = 4 {divided each side by 3}

the number is 4

For more help from me, visit: www.algebrahouse.com



Answer by ikleyn(52781)

.
You can solve it MENTALLY, without using equation/equations.


Your first step is to divide 42 by 6.


Do it, and then think what is your second step.